7. Groups of people aged​ 15-65 are randomly selected and arranged in groups of six. In the accompanying​ table, the random variable x is the number in the group who say that their family​ and/or partner contribute most to their happiness​ (based on a​ survey).

x   P(x)
0   0+
1   0.013
2   0.036
3   0.186
4   0.176
5   0.205
6   0.384

If the table is a probability​ distribution, what is its​ mean? Select the correct choice below and fill in any answer boxes within your choice.

Its mean is___

(Round to one decimal place as​ needed.)

The table is not a probability distrution
Answer #1

If  \sumP(X) = 1 then given distribution is probability distribution.

Our \sumP(X) = 0+0.013+0.036+0.186+0.176+0.205+0.384 = 1

So the given table is probability distribution table

The mean = \mu = \sumX * P(X)

=0*0+1*0.013+2*0.036+3*0.186+4*0.176+5*0.205+6*0.384

Mean = 4.7